OpenStudy (anonymous):

Helppppppp. Solve the system: 2x + 4y = 36 10y = 5 <3

OpenStudy (anonymous):

2x + 4y = 36 (divide the entire equation by 2) x + 2y = 18 (subtract 2y from both sides) x = 18 - 2y

OpenStudy (anonymous):

hot dang ur good at math

OpenStudy (anonymous):

You fn rock. :3

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

hershey look at mine and help :(

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

post link

OpenStudy (anonymous):

6. The transformation y=f(x) +k does what to the graph?

OpenStudy (lgbasallote):

what a fast typer hershey is...and fast net